# Born a Transformer – Always a Transformer?

### v2

@@ 4.2 In-Context Copying in Natural Settings as a Test Bed
In Section 4.1, forward copying showed near-perfect performance irrespective of uniqueness (UF and NF). We now ask whether the theoretically predicted asymmetry between unique and non-unique copying still impacts LLM behavior, by having models reproduce longer, naturalistic but semantically neutral text. We here do not focus on evaluating exact copying accuracy, instead focusing on understanding the nature of LLM mistakes. While entirely incorrect outputs are easily detectable, subtle differences in nearly accurate copies represent the most challenging errors for users to identify. Therefore, we focus exclusively on cases where the generated output is within 75% of the length of the expected output, targeting scenarios where the model clearly understands the copying task and produces plausible copies of the input. To confirm that our prompting setup is valid, we evaluate the models on copying exact spans from recent arXiv papers (April 2025), finding perfect performance.
Copying Lorem Ipsum Paragraphs.
We design our copying benchmark to suppress semantic cues that could help the model break ambiguous copy chains. Specifically, each input consists of randomized Lorem-Ipsum–style paragraphs—content that is unlikely to have appeared verbatim in the training corpus. Each paragraph has close to 500 tokens (as experiments in Section 4.1 exhibited perfect performance on lower string lengths).
@@
We categorize the transition indices as unambiguous or ambiguous based on nature of the token at that index. Since an induction head would suffice to correctly predict the subsequent token following an unambiguous token, we hypothesized that transition indices would never be unambiguous. Conversely, ambiguous tokens were expected to be the source of where all of the misaligned error span began, as it has similarities to the failure modes expected in NF (Non-unique forward copying). Empirical results (Figure 4) confirm this hypothesis. Across all model variants tested (sizes, families, completion as well instruction tuned models), a consistent pattern emerges, the transition indices almost unilaterally always are ambiguous and cause the ‘glitches’ we see in the final copied output. While these glitches appear across models, their exact nature appears to be random. We did not find any pattern except for the fact that they were tied to ambiguous tokens.
Git Commit History: Direction-Aware Copying.
@@
While the Lorem Ipsum setup shows how glitches in forward copying might cause reliability issues, it is not clear how similar glitches in backwards copying impact real world tasks. We use a compact Git history manipulation task to exemplify when the right-over-left asymmetry seen in Section 4.1 surfaces in practice. Given a truncated git log (newest→oldest commit hashes), the model must list commits that need revertinggit revert(forward order) or must be cherry-pickedgit cherry-pick(reverse order) onto a release branch. Llama-3 models reproduce the forward list almost perfectly but accuracy drops when the order flips (Figure 5) – mirroring the asymmetry seen in the synthetic experiments (Section 4.1). Qwen models are not able to do the task at all, hence we exclude them from this analysis. The reversal presented here does not happen here at the token level, instead at commit order level and is thus in principle more similar to a generalized copying setup, however the theoretical results and expectations transfer to the generalized scheme (See Appendix for a detailed discussion A.2.1).
@@
4.3 Fine-tuning Eliminates the Asymmetry, But Not the Glitches
@@
Our experiments have uncovered three failure modes: (i) a directional bias in pretrained models, (ii) poor performance on certain retrieval tasks that are provably length-generalizable, and (iii) sporadic “glitches” on copying tasks that are not length-generalizable. Each of these could stem either from architectural limitations or from inductive biases imparted during pretraining. To disentangle the two, we carry out task-specific supervised fine-tuning and see if we can get rid of the failure mode. Since we wish to know whether the model has truly learned a length-generalizable solution, we evaluate on test sequences twice as long as those seen in training, following Huang et al. (2025). A fully correct model should therefore achieve $100\%$ accuracy on this test set; even a single error then flags a reasoning flaw that could undermine reliability in real-world applications, an evaluation scheme similar to that of Liu et al. (2024a).
@@
Figure 6: On the out-of-distribution dataset, tasks belonging to C-Rasp[Pos] (UL, UR, UF, UB, NLFirst, NRFirst) retain perfect accuracy, whereas NLLast NRLast, NF and NB, which are provably not in C-Rasp[Pos], suffer a drop in performance. The left-right asymmetry within task setups largely disappears after fine-tuning.
@@
Experimental Setup. We fine–tuned a pre–trained language model —GPT–2 1.5B (APE) [^6] https://huggingface.co/karpathy/gpt2_1558M_final4_hf— on each retrieval task plus unique forward and backward copying. Fine–tuning uses the standard next–token objective but masks the loss to the answer span only, mimicking inference, i.e., for retrieval tasks, just the answer token, and, for the copying tasks, all tokens after the EOS delimiter in the input, indicating the end of the input sequence. To ensure that APE models can generalize on longer positions than those seen during training, we apply the positional offset trick of Huang et al. (2025): for every training input a random constant $\Delta\in[0,200-\ell_{min}]$ is added to all position indices, ensuring that all the positional embeddings are trained. All models see training examples whose input lengths are sampled uniformly from $[\ell_{\min},100]$, where $\ell_{\min}$ is the length of the shortest well–formed instance of the task. Details of hyperparameters and datasets are in Appendix D.
@@
Our testbed setup for all of our tasks closely follows the setup of Bhattamishra et al. (2020); Huang et al. (2025), with an in-distribution bin (i.e., lengths $[\ell_{min},100]$) and an out-of-distribution evaluation bin (i.e., lengths $[101,200]$). Similar to their setup, we report accuracy on our test sets only on seeds where training loss converges to 0. For our out-of-distribution evaluation bin on Retrieval tasks, we enforce that all instances of the query token $q$ appear exclusively in the first half of the context $\mathcal{C}$—increasing the distance between relevant tokens to which the model needs to attend, thereby rigorously testing the model’s ability to generalize to longer contexts. In such settings, superficial heuristics should be insufficient, and thus we expect that only models with robust length generalization can perform reliably. We find that C-Rasp[pos] perfectly tracks length generalization results in our fine-tuning experiments: Every task which belongs to C-Rasp[pos], and is thus predicted to length-generalize, does so, whereas tasks not in C-Rasp[pos] do not length-generalize. Additionally, the asymmetry seen during pretraining also goes away with fine-tuning, confirming the fact that it was an artifact of pretraining. We next seek to understand what leads to the development of this asymmetry by analyzing induction and anti-induction circuits (defined in Section 4.4) both in pretrained models, where this asymmetry exists, and in our fine-tuned models, where it doesn’t.
4.4 Source of the Asymmetry: A Mechanistic Perspective
@@
We hypothesize that the source of asymmetry in pretrained models lies in the difference between the presence and strength of induction and anti-induction circuits as described in Section 2. We posit that, in pretrained models, anti-induction heads are less common, but that fine-tuning can boost these, removing the asymmetry. We use Unique copying (UF and UB) setups to evaluate our hypothesis.
@@
Figure 7: (a) Top-10-percent attention heads ranked by their attention score, i.e. the mean attention weight that target-string tokens allocate to either the same source token (anti-induction) or the next source token (induction). Dashed lines:pretrained checkpoints; solid lines: fine-tuned. Shaded bands cover source lengths 3–10. Fine-tuning for forward copying (UF) strengthens induction heads, fine-tuning for backwards copying (UB) boosts anti-induction heads—eliminating the pretraining imbalance. (b) Patching experiment that ablates each family of heads while measuring task accuracy. For UF, removing induction heads (red $\rightarrow$ green) is fatal, whereas ablating anti-induction heads is inconsequential (red $\rightarrow$ blue); the roles reverse for UB. This holds across models & sequence lengths confirming the causal status of the two circuits.
Induction and anti-induction heads are causal in both fine-tuned and pretrained models. We conduct a patching experiment to confirm that removal of induction heads hurts performance in the UF task, while removal of anti-induction heads hurts performance in the UB task. Indeed, both in fine-tuned and pretrained models, removing induction heads is catastrophic for UF and has a negligible effect in UB, and removing anti-induction heads cripples backward copying while leaving forward copying intact. On fine-tuned checkpoints, the effect is amplified to the extreme: removing relevant heads drops the accuracy to zero, while removing another type of heads has no effect (Figure 77(b)). Thus, induction and anti-induction heads have a causal effect on model performance (Details of patching experiment in Appendix E). Additionally, the induction and anti-induction heads get amplified via fine-tuning (Figure 77(a)) Fine-tuning strengthens the heads that matter: induction heads gain attention weight for forward copy, anti-induction heads for backward copy.
Takeaway: The asymmetry in pretrained transformers can be explained by the difference in strength and prevalence of induction vs. anti-induction heads. Fine-tuning can be used to amplify either of these heads and thus remove the asymmetry.
@@ 6 Conclusion
We asked whether large-scale pretraining can erase the transformer’s native length-generalization limits, using a foundational family of retrieval and copying tasks as testbed. We find that pretraining boosts certain abilities, but transformers remain bound by the same length-generalization limits they were “born” with: Pretraining amplifies induction circuits—making right-/forward-oriented retrieval and copying notably easier—but does not overcome the difficulty inherent to non-unique retrieval and copying. Fine-tuning can rebalance induction and anti-induction heads and restore full, theory-aligned generalization, but only by explicitly teaching the model its missing capacities. Our results thus suggest that, while large-scale pretraining brings remarkable capabilities, it cannot fundamentally rewrite the architecture’s core inductive biases.
@@ Acknowledgements
Funded in part by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 232722074 – SFB 1102 “Information Density and Linguistic Encoding”; Project-ID 471607914 – GRK 2853/1 “Neuroexplicit Models of Language, Vision, and Action”; and Project-ID 389792660 – TRR 248 “Foundations of Perspicuous Software Systems”. We would like to thank Mark Rofin, Blerta Veseli, Brian DuSell, Ekaterina Shutova, Kayo Yin, Xinting Huang, Yifan Wang for discussions and feedback on the draft.
@@
YV contributed to Sections 4.1, 4.2, 4.3, 4.4, and paper refining. MJ contributed to Sections 4.1, 4.2, 4.3, scaling experiments, paper refinement and the Appendix. YS contributed to Sections 4.1, 4.2, the Appendix, scaling experiments, paper drafting and refining. SB contributed to Section 4.4. VD and EP provided feedback, supervision, and refined the paper. MH supervised the project and contributed to Section 3, paper drafting, and refining.
@@ A.1 C-Rasp[pos] as a Framework for Predicting Length Generalization
Let $\Phi$ be a set of unary relations ϕ:{,}:italic-$\phi:\mathbb{N}\rightarrow\{\bot,\top\}$ where each $\phi$ satisfies
@@
  • (Depends only on Distance) $\psi(i,j)$ only depends on $i-j$
@@
[…]
@@
The most important example for a function $\psi\in\Psi$ is the “predecessor” predicate checking if $j=i-1$:
@@
We will follow Huang et al. [2025], Yang and Chiang [2024] in using, as syntactic sugar, $\exists[j\leq i,\psi(i,j)]P(j)$ as a shorthand for $1\leq\#[j\leq i,\psi(i,j)]P(j)$, and #$\textsc{{\#}}\left[j\leq i\right]\;P(j)$ for #$\textsc{{\#}}\left[j\leq i,\top\right]\;P(j)$.
In this paper, we are interested in tasks that, given a prefix, produce a token: either the overall output symbol in a retrieval task, or the next output symbol when copying a string. To formalize this, we assume that every program contains Boolean-valued operations named $NEXT_{a}(i)$ for each $a\in\Sigma$ such that $NEXT_{a}(i)$ holds if and only if the correct next symbol is $a$. Thus, we say that a task is expressible in C-Rasp[pos] if there is a program defining $NEXT_{a}(i)$ for each $a\in\Sigma$ such that, for each $i$ in the span of the output (a single position for a retrieval task, the full output string for a copying task), $NEXT_{a}(i)$ holds (i.e, returns “true”) at the final position $i$ if and only if the correct next symbol is $a$. Generation stops when none of these operations return “true”.

### v3

@@ 4.2 Testing Uniqueness Bias and Directional Bias in Natural Settings
In Section 4.1, forward copying showed near-perfect performance irrespective of uniqueness (UF and NF). We now ask whether the theoretically predicted asymmetry between unique and non-unique copying still impacts LLM behavior, by having models reproduce longer, naturalistic but semantically neutral text. We here do not focus on evaluating exact copying accuracy, instead focusing on understanding the nature of LLM mistakes.
@@
While entirely incorrect outputs are easily detectable, subtle differences in nearly accurate copies represent the most challenging errors for users to identify. Therefore, we focus exclusively on the cases where the generated output is within 75% of the length of the expected output, targeting scenarios where the model clearly understands the copying task and produces plausible copies of the input. To confirm that our prompting setup is valid, we evaluate the models on copying exact spans from recent arXiv papers (April 2025), finding perfect performance.
We design our copying benchmark to suppress semantic cues that could help the model break ambiguous copy chains. Specifically, each input consists of a randomized Lorem‑Ipsum–style paragraph. Such content is unlikely to have appeared verbatim in the training corpus. Each paragraph has close to 500 tokens. See Appendix C.5.1 for results at longer lengths (upto 5k tokens).
@@
We categorize the transition indices as unambiguous or ambiguous based on the nature of the token at that index. Since an induction head would suffice to correctly predict the subsequent token following an unambiguous token, we hypothesized that transition indices would never be unambiguous. Conversely, ambiguous tokens were expected to be the source of where all of the misaligned error span began, as it has similarities to the failure modes expected in NF (Non-unique forward copying). Empirical results (Figure 4) confirm this hypothesis. Across all model variants tested (sizes, families, completion as well instruction tuned models), a consistent pattern emerges, the transition indices almost unilaterally always are ambiguous and cause the ‘glitches’ we see in the final copied output. While these glitches appear across models, their exact nature appears to be random. We did not find any pattern except for the fact that they were tied to ambiguous tokens.
Testing Directional Bias in Git Commit History: Direction‑Aware Copying.
While the Lorem Ipsum setup shows how glitches in forward copying might cause reliability issues, we now show how failures in backward copying impact real world tasks. We use a compact Git history manipulation task to exemplify when the right-over-left asymmetry seen in Section 4.1 surfaces in practice. Given a truncated git log (newest→oldest commit hashes), the model must list commits that need revertinggit revert(forward order) or must be cherry‑pickedgit cherry-pick(reverse order) onto a release branch.
@@
Llama‑3 models reproduce the forward list almost perfectly but accuracy drops when the order flips (Figure 5) – mirroring the asymmetry seen in the synthetic experiments (Section 4.1). Qwen models are not able to do the task at all, hence we exclude them from this analysis. The reversal presented here does not happen here at the token level, instead at the commit order level and is thus in principle more similar to a generalized copying setup. However, the theoretical results and expectations transfer to the generalized scheme (See Appendix for a detailed discussion A.2.1).
@@
4.3 Fine-tuning Eliminates the Directional Bias, But Not the Uniqueness Bias
@@
Our experiments have uncovered three failure modes: (i) a Directional Bias in pretrained models, (ii) poor performance on certain retrieval tasks that are provably length-generalizable, and (iii) the Uniqueness Bias: sporadic “glitches” on copying tasks that are not length-generalizable. Each of these could stem either from architectural limitations or from inductive biases imparted during pretraining. To disentangle the two, we carry out task-specific supervised fine-tuning and see which failures are removed. Since we wish to know whether the model has truly learned a length-generalizable solution, we evaluate on test sequences twice as long as those seen in training, following Huang et al. (2025). A fully correct model should therefore achieve $100\%$ accuracy on this test set; even a single error then flags a reasoning flaw that could undermine reliability in real-world applications, an evaluation scheme similar to that of Liu et al. (2024a).
@@
Figure 6: On the out-of-distribution dataset, tasks belonging to C-Rasp[Pos] (UL, UR, UF, UB, NLFirst, NRFirst) retain perfect accuracy, whereas NLLast, NRLast, NF, and NB, which are provably not in C-Rasp[Pos], suffer a drop in performance. The Directional Bias within task setups is largely absent from the fine-tuned models.
@@
Experimental Setup. We fine-tune a pretrained language model —GPT–2 1.5B (APE) [^7] https://huggingface.co/karpathy/gpt2_1558M_final4_hf— on each retrieval task plus unique forward and backward copying. Fine-tuning uses the standard next-token objective but masks the loss to the answer span only, mimicking inference, i.e., for retrieval tasks, just the answer token, and, for the copying tasks, all tokens after the EOS delimiter in the input, indicating the end of the input sequence. To ensure that APE models can generalize on longer positions than those seen during training, we apply the positional offset trick of Huang et al. (2025): for every training input a random constant $\Delta\in[0,200-\ell_{min}]$ is added to all position indices, ensuring that all the positional embeddings are trained. All models see training examples whose input lengths are sampled uniformly from $\ell_{\min},100]$, where $\ell_{\min}$ is the length of the shortest well-formed instance of the task. We confirmed that, prior to this fine-tuning, the base model’s performance on each task was effectively zero. Details of hyperparameters and datasets are in Appendix D.
@@
Our testbed setup for all of our tasks closely follows the setup of Bhattamishra et al. (2020); Huang et al. (2025), with an in-distribution bin (i.e., lengths $\ell_{min},100]$) and an out-of-distribution evaluation bin (i.e., lengths $[101,200]$). Similar to their setup, we report accuracy on our test sets only on seeds where training loss converges to 0. For our out-of-distribution evaluation bin on Retrieval tasks, we enforce that all instances of the query token $q$ appear exclusively in the first half of the context $\mathcal{C}$. This way distance between the relevant tokens to which the model needs to attend increases, thereby rigorously testing the model’s ability to generalize to longer contexts. In such settings, superficial heuristics should be insufficient, and thus we expect that only models with robust length generalization can perform reliably. We find that C-Rasp[pos] perfectly tracks length generalization results in our fine-tuning experiments: Every task which belongs to C-Rasp[pos], and is thus predicted to length-generalize, does so, whereas tasks not in C-Rasp[pos] do not length-generalize. Additionally, the Directional Bias of pretrained models disappears with fine-tuning, confirming the fact that it was an artifact of pretraining. We next seek to understand what leads to the development of this asymmetry by analyzing induction and anti-induction circuits (defined in Section 4.4) both in pretrained models, where this asymmetry exists, and in our fine-tuned models, where it doesn’t.
4.4 Source of the Directional Bias: A Mechanistic Perspective
@@
We hypothesize that the source of directional bias in pretrained models lies in the difference between the presence and strength of induction and anti-induction circuits as described in Section 2. We posit that, in pretrained models, anti-induction heads are less common, but that fine-tuning can boost these, removing the asymmetry. We use Unique copying (UF and UB) setups to evaluate our hypothesis.
@@
Figure 7: (a) Top-10-percent attention heads ranked by their attention score, i.e. the mean attention weight that target-string tokens allocate to either the same source token (anti-induction) or the next source token (induction). Dashed lines: pretrained checkpoints; solid lines: fine-tuned. Shaded bands cover source lengths 3–10. Fine-tuning eliminates the pretraining imbalance; for forward copying (UF) fine-tuning strengthens induction heads, for backwards copying (UB) fine-tuning boosts anti-induction heads. (b) Patching experiment that ablates each family of heads while measuring task accuracy. For UF, removing induction heads (red $\rightarrow$ green) is fatal, whereas ablating anti-induction heads is inconsequential (red $\rightarrow$ blue); the roles reverse for UB. This holds across models & sequence lengths confirming the causal status of the two circuits.
Induction and anti-induction heads are causal in both fine-tuned and pretrained models. We conduct a patching experiment to confirm that removal of induction heads hurts performance in the UF task, while removal of anti-induction heads hurts performance in the UB task. Indeed, both in fine-tuned and pretrained models, removing induction heads is catastrophic for UF and has a negligible effect in UB, and removing anti-induction heads cripples backward copying while leaving forward copying intact. On fine-tuned checkpoints, the effect is amplified to the extreme: removing relevant heads drops the accuracy to zero, while removing another type of heads has no effect (Figure 77(b)). Thus, induction and anti-induction heads have a causal effect on model performance (Details of patching experiment in Appendix E). Additionally, the induction and anti-induction heads get amplified via fine-tuning (Figure 77(a)). Fine-tuning strengthens the heads that matter: induction heads gain attention weight for forward copy, anti-induction heads for backward copy.
Takeaway: The directional bias in pretrained transformers can be explained by the difference in strength and prevalence of induction vs. anti-induction heads. Fine-tuning can be used to amplify either of these heads and thus remove the asymmetry.
@@ 6 Conclusion
We asked whether large‐scale pretraining can erase the transformer’s native length‐generalization limits, using a foundational family of retrieval and copying tasks as testbed. We find that pretraining boosts certain abilities, but transformers remain bound by the same length-generalization limits they were “born” with: Pretraining amplifies induction circuits, making right-/forward-oriented retrieval and copying notably easier, but does not overcome the difficulty inherent to non-unique retrieval and copying. Fine-tuning can rebalance induction and anti-induction heads and restore full, theory-aligned generalization, but only by explicitly teaching the model its missing capacities. Our results thus suggest that, while large-scale pretraining brings remarkable capabilities, it cannot fundamentally rewrite the architecture’s core inductive biases.
@@ Acknowledgements
We thank anonymous reviewers for their encouraging and constructive feedback. This research is funded in part by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 232722074 – SFB 1102 “Information Density and Linguistic Encoding”; Project-ID 471607914 – GRK 2853/1 “Neuroexplicit Models of Language, Vision, and Action”; and Project-ID 389792660 – TRR 248 “Foundations of Perspicuous Software Systems”. We would like to thank Mark Rofin, Blerta Veseli, Brian DuSell, Ekaterina Shutova, Kayo Yin, Xinting Huang, Yifan Wang for discussions and feedback on the draft.
@@
MJ contributed to Sections 4.1, 4.2, 4.3, scaling experiments, paper refinement and the Appendix. YV contributed to Sections 4.1, 4.2, 4.3, 4.4, and paper refining. YS contributed to Sections 4.1, 4.2, the Appendix, scaling experiments, paper drafting and refining. SB contributed to Section 4.4. VD and EP provided feedback, supervision, and refined the paper. MH supervised the project and contributed to Section 3, paper drafting, and refining.
@@ A.1 C-Rasp[pos] as a Framework for Predicting Length Generalization
Let $\Phi$ be a set of unary relations $\phi:\mathbb{N}\rightarrow\{\bot,\top\}$ where each $\phi$ satisfies
@@
  • (Depends only on Distance) $\psi(i,j)$ only depends on iji-j
@@
A C-Rasp[pos] program $P$ is defined as a sequence $P_{1},\ldots,P_{k}$ of operations, each of the following types:
@@
The most important example for a function $\psi\in\Psi$ is the “predecessor” predicate checking if j=i1j=i-1:
@@
We will follow Huang et al. [2025], Yang and Chiang [2024] in using, as syntactic sugar, $\exists[j\leq i,\psi(i,j)]P(j)$ as a shorthand for $1\leq\#[j\leq i,\psi(i,j)]P(j)$, and $\textsc{{\#}}\left[j\leq i\right]\;P(j)$ for $\textsc{{\#}}\left[j\leq i,\top\right]\;P(j)$.
In this paper, we are interested in tasks that, given a prefix, produce a token: either the overall output symbol in a retrieval task, or the next output symbol when copying a string. To formalize this, we assume that every program contains Boolean-valued operations named $\text{NEXT}_{a}(i)$ for each $a\in\Sigma$ such that $\text{NEXT}_{a}(i)$ holds if and only if the correct next symbol is $a$. Thus, we say that a task is expressible in C-Rasp[pos] if there is a program defining $\text{NEXT}_{a}(i)$ for each $a\in\Sigma$ such that, for each $i$ in the span of the output (a single position for a retrieval task, the full output string for a copying task), $\text{NEXT}_{a}(i)$ holds (i.e, returns “true”) at the final position $i$ if and only if the correct next symbol is $a$. Generation stops when none of these operations return “true”.